OpenStudy (anonymous):

If f(x) = x + 3, find the value of x if f(x) = 9.

OpenStudy (anonymous):

Set f(x) equal to 9. 9 = x + 3 Can you solve for x?

OpenStudy (anonymous):

Plug in 9 where the x is. 9+3 =12

OpenStudy (anonymous):

f(x)=x+3 since f(x)=9 so, 9=x+3 x=9-3 x=6

OpenStudy (anonymous):

@IStutts You are incrorect. It says FIND x if f(x) = 9. Not find f(9).

OpenStudy (anonymous):

so its 12

OpenStudy (anonymous):

No. I just said that @IStutts was incorrect.

OpenStudy (anonymous):

so 6

OpenStudy (anonymous):

yup you got the right answer